• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Sirca

    Log entry time 17:35:04 on April 7,2001

    Entry number 60009

    keyword=Target probably misaligned

    When getting ready for taking data, we inserted the BeO target and put
    5muA on it. The beam spot as seen on the TV screen was displaced by
    about -4mm VER, and -4mm HOR (not sure about the sign of HOR)
    and was unusually diffuse. The "B" BPM was reading a perfect (0,0)
    and seems trustworthy. (Run #3378 was running with the BeO in.)


    Consulted cryo-target experts. There is a large probability that the target
    ladder is misaligned again, and that the motion mechanism might break
    again soon. The consensus was to move the LH2 into position and resume
    production running during the weekend. Even if the misalignment is
    genuine and the LH2 target does not land exactly in the designated space,
    an offset of 4mm should not render the data useless -- will estimate
    over the weekend whether this is true, based on the analysis of different
    angular settings.